### v1

# Robo-Instruct: Simulator-Augmented Instruction Alignment For Finetuning CodeLLMs

Abstract
Large language models (LLMs) have shown great promise at generating robot programs from natural language given domain-specific robot application programming interfaces (APIs). However, the performance gap between proprietary LLMs and smaller open-weight LLMs remains wide. This raises a question: Can we fine-tune smaller open-weight LLMs for generating domain-specific robot programs to close the performance gap with proprietary LLMs? While Self-Instruct is a promising solution by generating a diverse set of training data, it cannot verify the correctness of these programs. In contrast, a robot simulator with a well-defined world can identify execution errors but limits the diversity of programs that it can verify. In this work, we introduce Robo-Instruct, which brings the best of both worlds — it promotes the diversity of Self-Instruct, while providing correctness of simulator-based checking. Robo-Instruct introduces RoboSim to synthesize a consistent world state on the fly by inferring properties relevant to the program being checked, and simulating actions accordingly. Furthermore, the instructions and programs generated by Self-Instruct may be subtly inconsistent — such as the program missing a step implied by the instruction. Robo-Instruct further addresses this with InstAlign, an instruction-program alignment procedure that revises the task instruction to reflect actual results of the generated program. Given a few seed task descriptions and the robot APIs, Robo-Instruct is capable of generating a training dataset using only a small open-weight model. This dataset is then used to fine-tune small open-weight language models, enabling them to even exceed the performance of several proprietary LLMs including GPT-3.5-Turbo and Gemini-Pro.
Figure 1: High-Level Overview of Robo-Instruct. This figure also illustrates an example of an invalid Self-Instruct-generated instruction and program, as well as pass@1 results of different LLMs on RoboEval.
1 Introduction
Large language models (LLMs) have demonstrated great promise at generating robot programs from natural language instructions [10, 12, 3, 18, 39, 17, 31, 11]. For example, consider an instruction for a service mobile robot: "Check how many conference rooms have no markers." The robot may be equipped with a domain-specific robot application programming interface (API) that includes skills such as go_to(location) for navigation and is_in_room(object) for perception. Since such domain-specific APIs do not exist in the training dataset of general-purpose LLMs, in-context learning (ICL) via few-shot examples is often employed to describe and use such APIs for performing few-shot inference. However, there is a significant performance gap [10] in the correctness of programs generated by ICL for large proprietary models and smaller open-weight models that can be deployed locally on robots. This raises a question: can we fine-tune small open-weight LLMs for generating domain-specific robot programs to close the performance gap with proprietary LLMs?
Since training datasets of the domain-specific robot programs are often unavailable, Self-Instruct might seem like a promising solution [36, 29]. Consider the setting of generating programs for service mobile robots that can perceive objects, navigate to various locations, manipulate items, and communicate with humans. By formulating these robot skills into APIs, we can create a few seed task examples demonstrating their use case and employ Self-Instruct to generate a diverse set of instruction-program pairs as training data, as illustrated in Fig. 1. However, using Self-Instruct naïvely may generate infeasible instructions—e.g., asking the robot to pick up multiple objects at once when it cannot due to physical constraints. They can also violate domain-specific constraints. For example, in Fig. 1, after line 2 confirms the absence of a key at the current location, line 3 erroneously attempts to pick up a key. Further, these instructions may not align with the generated programs, even if these programs are valid. For example, Fig. 1 shows an example instruction directing the robot to verbally ask in each room if a key exists, whereas the program instructs the robot to visually check in each room. Finally, the generated programs may have execution errors. These challenges may appear to be solvable using a simulator, but a simulator needs an initial world state to check against programs. A simulator using a hand-curated world state will end up rejecting the wide diversity of programs generated by Self-Instruct, even if they are executable, just because the world state did not capture some aspect relevant to them (e.g., the presence of a “key”).
This work introduces Robo-Instruct, a new framework based on Self-Instruct, to address these issues and improve the performance of small open-weight language models for generating domain-specific robot programs. As shown in Fig. 1, Robo-Instruct introduces two novel components: (1) RoboSim, a task-agnostic simulator that encodes domain-specific constraints and validates robot programs generated from Self-Instruct. Critically, RoboSim dynamically synthesizes a consistent world state starting from arbitrary programs. (2) InstAlign, an instruction-program alignment procedure that revises the generated instructions to better reflect the intent of the generated programs. Robo-Instruct also employs a rejection-sampling mechanism that rejects invalid programs detected by RoboSim and queries Self-Instruct for a new program corresponding to the same generated instruction.
We validate Robo-Instruct by fine-tuning Codellama-Python-7B [30] and evaluate on RoboEval, a domain-specific code generation benchmark for service mobile robots. We show that Robo-Instruct is capable of improving the performance of the Codellama model by using only a small open-weight model to generate the training dataset. Compared to the base Codellama-Python-7B model without fine-tuning, our Robo-Instruct fine-tuned models outperform by $28.75\%$ in average pass@1 scores; and, compared to Self-Instruct fine-tuned model, our model outperform by $13.75\%$.; and the best pass@1 of Robo-Instruct fine-tuned model achieves a 68.75% match, surpassing the performance of the proprietary GPT-3.5-Turbo and Gemini-1.0-Pro.
Contributions
Our main contributions are as follows:  [^1] Website coming soon: https://amrl.cs.utexas.edu/robo-instruct/
  1. 1.
    We introduce Robo-Instruct, a new framework for improving the code generation performance of small open-weight language models for domain-specific robot programs. This framework introduces two novel components, RoboSim and InstAlign.
  2. 2.
    We introduce a dynamic world synthesis and evaluation process for generating relevant world states for automated code checking for diverse, arbitrary tasks in RoboSim.
  3. 3.
    We introduce InstAlign, an instruction alignment procedure to refine instruction-code pairs to improve alignment between instructions and code generated by Self-Instruct.
  4. 4.
    We fine-tune a small open-weight model, Codellama-Python-7B [30], using Robo-Instruct, and improve its performance to outperform several CodeLLMs, including Deepseek-Coder-33B [8], and Starcoder2-15B [21] and two proprietary LLMs, GPT-3.5-Turbo [27] and Gemini-1.0-Pro [33] on the RoboEval benchmark.
2 Robo-Instruct
In this section, we present how Robo-Instruct generates training datasets of domain-specific robot programs. Alg. 1 shows a broad overview of the framework. To add an entry in the training dataset, Self-Instruct first generates an instruction-program pair, $(\mathcal{I},\mathcal{P})$, from the robot APIs and seed tasks, shown in Appendix A.4. Then, RoboSim dynamically synthesizes a consistent world state on the fly as it executes and validates $\mathcal{P}$. If $\mathcal{P}$ is invalid, Robo-Instruct employs a rejection-sampling method, which generates a new program $\mathcal{P}$ given the same $\mathcal{I}$ and evaluates the new $\mathcal{P}$ again. This process repeats until $\mathcal{P}$ becomes valid or a predefined maximum resampling limit is reached. If the limit is reached, the instruction might be invalid given the domain-specific APIs or too complex to generate a program, so the instruction-program pair is discarded. Finally, if $\mathcal{P}$ is valid, InstAlign takes in $(\mathcal{I},\mathcal{P})$ to revise $\mathcal{I}$ to better reflect the intent of $\mathcal{P}$ and the aligned instruction and program is saved to the training dataset. In the following subsections, we elaborate on the specific design of each component.
Algorithm 1 Robo-Instruct: Instruction-Program Generation
[…]
1:  Initialize: $\mathcal{D}=\emptyset$triangleright Training dataset
2:  Initialize: $N$triangleright Training dataset size
3:  Initialize: $m$triangleright Maximum resampling limit
4:  while $\text{len}(\mathcal{D})<N$ do
5:     $\mathcal{I}\leftarrow\textsc{Self-Instruct}_{inst}(\mathcal{S})$
6:     $\mathcal{P}\leftarrow\textsc{Self-Instruct}_{code}(\mathcal{S},\mathcal{I})$
7:     for $i=1$ to $m$ do
[…]
10:           $\mathcal{P}\leftarrow\textsc{Self-Instruct}_{code}(\mathcal{S},\mathcal{I})$triangleright Rejection-sampling
11:        else
12:           $\mathcal{I}_{\text{aligned}}\leftarrow\textsc{InstAlign}(\mathcal{S},\mathcal{I},\mathcal{P})$triangleright Align instruction with the program
13:           $\mathcal{D}\leftarrow(\mathcal{I}_{\text{aligned}},\mathcal{P})$
14:           break
15:        end if
16:     end for
17:  end while
18:  return  $\mathcal{D}$
2.1 RoboSim: A Task-Agnostic Simulator For Domain-Specific Programs
Algorithm 2 $\textsc{RoboSim}(\mathcal{P})$
0:  Program $\mathcal{P}$triangleright Generated program
1:  Initialize: Set $\mathcal{A}$triangleright A set of domain-specific robot APIs
2:  Initialize: $k$triangleright Number of evaluation iterations
3:  Initialize: $\mathcal{W}_{\text{init}}$triangleright An initial world state with or without predefined information
4:  $\mathcal{P}_{\text{trans}}\leftarrow\textsc{Translate}(\mathcal{P},\mathcal{A},\textsc{DynamicEval})$triangleright Replace each API call with DynamicEval
5:  for $i=1$ to $k$ do $\triangleright$ Then, evaluate $\mathcal{P}$ $k$ times to catch program errors
6:   try:
7:    $\mathcal{W}\leftarrow\mathcal{W}_{\text{init}}$triangleright Initialize a new world state
8:    exec$(\mathcal{P}_{\text{trans}},\mathcal{W})$
9:   catch:
10:    return False
11:  end for
12:  return  True $\triangleright$ Return True if all program executions are successful
[…]
2:  for $l\in p$ do $\triangleright$ Loop through every literal in the precondition
3:   if $\textsc{CheckDefined}(\mathcal{W},l)==\text{undefined}$ then
4:    $W\leftarrow\textsc{GrowWorld}(l,\mathcal{W})$triangleright Instantiate the literal and grow $\mathcal{W}$ to include it
5:   end if
6:  end for
7:  $\text{retval},\mathcal{W}\leftarrow\textsc{ExecUpdate}(\text{api\_fn},\text{params},\mathcal{W})$triangleright Execute api_fn and update $\mathcal{W}$
8:  return  retval, $\mathcal{W}$
We present a principled approach to design RoboSim for validating domain-specific robot programs. Alg. 2 illustrates the high-level algorithm used to assess the correctness of a robot program. RoboSim employs the concept of world state to simulate the robot actions directed by a program, ensuring consistent and reliable evaluation. A world state is a symbolic representation of the environment in which the robot operates, and it keeps track of the high-level changes in the robot state and the surrounding environment as the robot performs actions in order. For example, consider a program instruction that commands a robot to check if a pen is nearby. The world state queries the stored information about the surrounding environment, identifies all objects at the robot’s current location, and informs the program whether a pen is present.
However, since Self-Instruct generates arbitrary programs based on the provided APIs, RoboSim does not know what a plausible world state relevant to the program would be a priori — e.g., reasoning about the existence of a pen in the example program shown in Fig. 1. Thus, we equip RoboSim with the ability to synthesize the world state as more robot actions are performed. Our approach is inspired by angelic execution [4], which has previously been used for software verification of programs with partially defined library functions. In our case, instead of partially defined library functions, we have unknown plausible world states. RoboSim dynamically synthesizes and grows a world state based on domain-specific constraints (e.g., object permanence, robot skills, etc.) and the execution trace of the program, which allows it to infer a consistent and relevant world state.
Specifically, RoboSim modifies the program to replace all API calls with the DynamicEval function (Alg. 2 line 4) — when an API function is called during execution, the DynamicEval function is invoked instead.
DynamicEval makes an important extension to the formulation of STRIPS [7] to integrate with API functions. DynamicEval equips each API function with specific pre-conditions, effects, and return values. The pre-conditions are composed of literals tailored to the function’s requirements. For instance, the API function is_in_room("pen"), which determines if an object ‘pen’ is in the same room as the robot, uses two literals for its pre-condition: robot_at(X) and obj_at(X, "pen"). Generally, STRIPS assigns one of two possible values to each literal: True if the literal is defined, otherwise False. However, prior to program execution, DynamicEval is unaware of the program-relevant literals. Thus we assign a third value, undefined, to such unknown literals. Literals must thus be explicitly defined as either True or False, or they remain undefined if not specified.
Alg. 3 demonstrates how DynamicEval executes an API function and updates the world state. First, it calculates the precondition specified for the function. It then checks each literal in the precondition to see if it is defined. If a literal is undefined, DynamicEval invokes GrowWorld, a stochastic function that assigns a random truth value to the literal and updates the world state accordingly. Finally, DynamicEval proceeds to execute the API function using the current world state, retrieves the return values, and applies the function’s effects to update the world state.
Fig. 2 illustrates an example of RoboSim executing a generated program. Initially, RoboSim’s world state only specifies the robot’s current location, and whether a pen is in the same room as the robot remains undefined (line 2). Therefore, DynamicEval invokes GrowWorld to randomly determine a truth value for the obj_at(start_loc, "pen") literal, leading to two distinct execution paths depicted in light purple and blue. Subsequently, as additional API functions are called, more literals are introduced or updated in the world state to ensure consistent evaluations. In addition, a crucial aspect of RoboSim’s design is that a literal can also be set to an undefined value. In world state 2 on the purple execution path, after executing the pick("pen") function, the program is unable to determine if another pen exists in the current room, given the provided APIs. As a result, the world state updates the literal obj_at(start_loc, "pen") from True to Undefined.
Finally, due to the stochastic nature of DynamicEval, RoboSim must execute the generated program multiple times to validate the program. If all executions are successful, the program is deemed correct (Alg. 2 line 5-11).
Figure 2: Example of RoboSim executing a generated program and updating the world state. Initially, RoboSim begins with a world state that includes only the robot’s current location. As the program executes, two distinct execution paths emerge, depicted in light purple and blue. This figure demonstrates how the world state is updated along each execution path.
2.2 InstAlign: Instruction-Program Alignment Procedure
Given that LLMs are extensively trained in code understanding [30], InstAlign is a procedure that prompts an LLM to revise $\mathcal{I}$ to better reflect the intent of $\mathcal{P}$. This procedure involves two steps: first, given $\mathcal{I}$ and $\mathcal{P}$, InstAlign leverages Chain-of-Thought reasoning [37] (CoT) to prompt an LLM to generate a revised instruction, $\mathcal{I}_{\text{revised}}$; second, InstAlign invokes the LLM again to determine whether $\mathcal{I}$ or $\mathcal{I}_{\text{revised}}$ is more aligned with $\mathcal{P}$’s intent and output the chosen instruction as $\mathcal{I}_{\text{aligned}}$.
To generate $\mathcal{I}_{\text{revised}}$, the prompt to the LLM comprises the robot API function definitions, $\mathcal{I}$, $\mathcal{P}$, and CoT instructions. The CoT asks the LLM to perform the following three steps in order: 1. write down all the robot APIs used in the program; 2. examine these APIs and write down step by step what the program does; 3. combine all the information above to revise the robot instruction. Similarly, to determine $\mathcal{I}_{\text{aligned}}$, an LLM is prompted to think step by step about $\mathcal{P}$, $\mathcal{I}$ and $\mathcal{I}_{\text{revised}}$ to arrive at a conclusion. Detailed prompt is shown in Appendix A.6.
3 Analysis and Experiments
In this section, we investigate the following two research questions:
  1. 1.
    Is Robo-Instruct effective at generating training data to fine-tune a small language model for generating domain-specific robot programs?
  2. 2.
    How do RoboSim and InstAlign impact the effectiveness of Robo-Instruct?
We conduct our investigation by fine-tuning the Codellama-Python-7B model [30] on the synthetic dataset generated by Robo-Instruct and evaluate the fine-tuned model using RoboEval [10], a domain-specific code generation benchmark for service mobile robots. In the following subsections, we first provide a brief description of RoboEval. Then we present our experimental results addressing the two main research questions. Finally, we offer more analysis of RoboSim, InstAlign, and the synthetic dataset.
3.1 RoboEval: A Domain-Specific Robot Code Generation Benchmark
Figure 3: RoboEval APIs and benchmark task example.
RoboEval is a domain-specific code generation benchmark, featuring a suite of 16 tasks designed to evaluate the ability of LLMs to understand custom APIs and generate programs for service robots. In this domain, a service robot can perceive objects, navigate to various locations, manipulate items, and communicate with humans. Furthermore, the robot should be capable of basic commonsense reasoning and executing complex tasks that involve conditional and repetitive actions. To facilitate these capabilities, RoboEval defines a set of 8 API functions in Python as skill primitives. Fig. 3 illustrates these function signatures and definitions, alongside an example task instruction and its canonical solution from the benchmark. In addition, unlike other popular code generation benchmark tasks [6, 2, 16, 19, 14, 9], the order of the robot’s actions is crucial for successfully completing the specified tasks. For instance, in the task "bring me a marker from the classroom that does not have a whiteboard," the robot must check each classroom until it finds one without a whiteboard, whereas simply bringing back a marker is insufficient. Hence, RoboEval evaluates the generated program by executing it in a simulator to capture the action traces, which are subsequently validated for sequence correctness using temporal logic.
3.2 RQ1: Is Robo-Instruct Effective at Generating Training Data to Fine-Tune a Small Language Model for Generating Domain-Specific Robot Programs?
RoboEval pass@1
Fine-tune Model # Param $T=0$ $T=0.2$ Licensing
- GPT-4 - 83.75% 85.81% Proprietary
- GPT-3.5 - 67.5% 65.56% Proprietary
- Gemini-1.0-Pro - 60.00% 59.88% Proprietary
- Codellama-Python 7B 40.00% 39.31% Open
- Codellama-Python 34B 46.25% 48.25% Open
- Starcoder2 15B 62.5% 60.94% Open
- Deepseek-Coder 33B 53.75% 52.13% Open
- Llama3-Inst 8B 48.75% 48.38% Open
Self-Instruct Codellama-Python 7B 55.00% 52.69% Open
Robo-Instruct (ours) Codellama-Python 7B 68.75% 66.00% Open
Table 1: Pass@1 results of different LLMs on RoboEval computed with greedy decoding $T=0$ and nucleus sampling $T=0.2$.
Experiment Setup. We use the open-weight LLM, Llama3-8B-Inst, for Robo-Instruct. To generate a diverse dataset, we employ nucleus sampling for creating instruction-program pairs, setting the temperature $T=1$ and top $p=0.95$. The maximum resampling limit is capped at $3$ to accommodate instructions that initially produce invalid programs. For the LLM used in InstAlign, we empirically adjust the generation temperature to $T=0.3$ to optimize performance. Furthermore, we assess the edit similarity between token sequences of each instruction pair in the dataset [15], removing duplicates where the similarity score exceeds 0.6. We use the same setup to generate data via Self-Instruct. Instead of discarding invalid programs, Self-Instruct includes every generated instruction-program pair in the training dataset. Finally, we create two datasets with 5K instruction-program pairs each using Self-Instruct and Robo-Instruct respectively. These datasets are then used to fine-tune the Codellama-Python-7B model. The learning rate is set to be $-53e\text{-}$ 5 with a warmup ratio of $3\%$ and a constant lr scheduler. We employ the AdamW optimizer [20] with an effective batch size of 8, training each model for 5 epochs using a sequence length of 2048 tokens. We train all our models on a single H-100 GPU using unsloth [35].
Baselines. We divide our baseline models into 2 categories: 1) proprietary LLMs, including GPT4 [28], GPT3.5-Turbo [27], Gemino-Pro [33], and 2) open-weight LLMs, including Codellama-Python-7B [30], Codellama-Python-34B, Starcoder2-33B [21], Deepseek-Coder-33B [8], and Llama3-8B-Inst [1]. All the results are evaluated using RoboEval and reported in Tab. 1.
Tab. 1 presents the average pass@1 results for different LLMs on RoboEval, using two different temperature settings for generation: greedy decoding at a temperatures of $T=0$ and nucleus sampling at a temperature of $T=0.2$. The results show that Robo-Instruct-fine-tuned Codellama significantly improves upon the base Codellama-Python-7B and outperforms the Self-Instruct-fine-tuned variant. Notably, it surpasses all open-weight models, including larger ones like Codellama-Python-34B and Deepseek-Coder-33B. Additionally, although the training dataset was generated using Llama3-8B-Inst, which scores less than 50% pass@1 on RoboEval, our Robo-Instruct-fine-tuned model still achieves a significant improvement, scoring 68.75% under deterministic temperature settings for generation. Finally, compared to proprietary models, while our Robo-Instruct-fine-tuned model trails the more powerful GPT-4, it outperforms GPT-3.5-Turbo and Gemini-1.0-Pro in generating programs for service mobile robots. This result demonstrates the effectiveness of our approach in generating domain-specific robot program data for fine-tuning a small language model. It suggests that the fine-tuned model could potentially replace some proprietary models, providing a more cost-effective and private option for local deployment.
3.3 RQ2: How Do RoboSim and InstAlign Impact the Effectiveness of Robo-Instruct?
T=0 T=0.2 Invalid
Method pass@1 Improv. pass@1 Improv. Programs
Codellama-7B-Python 40.00% +0% 39.31% +0% 38.31%
Self-Instruct 55.00% +15.00% 52.69% +13.38% 20.94%
[…]
Table 2: Pass@1 results of different methods on RoboEval computed with greedy decoding $T=0$ and nucleus sampling $T=0.2$. The Invalid Programs column indicates the percentage of programs that result in execution errors when tested on RoboEval tasks.
[…]
3.4 Qualitative analysis of the generated program errors
Figure 4: Self-Instruct-Generated Program Errors: Examples 1 to 4 illustrate errors specific to the Python language, and Examples 5 to 8 highlight errors rooted in domain-specific constraints. [^3] Programs have been adapted to succinctly demonstrate the types of errors and fit within the figure.
We analyze invalid programs identified by RoboSim, categorizing the errors into two types: language-native errors and domain-specific constraint violations. Fig. 3 displays eight examples of these programs, with Examples 1 to 4 illustrating errors specific to the Python language, and Examples 5 to 8 highlighting errors rooted in domain-specific constraints. Language-native errors are generally straightforward, such as syntax errors, the use of undefined variables or functions, or improper use of provided APIs.
In contrast, errors related to domain-specific constraints tend to be more complex to detect. For instance, Example 5 illustrates the program incorrectly trying to pick up a watering can (line 3) after establishing that it is not present at the location (line 2). Similarly, Example 6 demonstrates an error where the program inappropriately asks Jack (line 5) after confirming his absence from the room (line 3). Example 7 illustrates a scenario in which RoboSim updates the world state by labeling "item storage room" as a location after executing the go_to command (line 2). Subsequently, the robot attempts to pick up this location (line 3), resulting in an error. Example 9 is the most intricate scenario where the world state in the living room is updated to include a toy after the robot places it there (line 7). When the robot returns to the living room for the second time (line 5), it does not place down what it holds (line 7). Hence, in the third room the robot visits (line 3), when it attempts to pick up a toy again (line 4), an error occurs because the robot can only carry one item at a time.
4 Related Work
4.1 LLMs for Robot Code Generation
LLMs have shown impressive capabilities in generating robot programs from natural language [17, 31, 11]. One popular approach uses LLMs to generate composable costmaps for robots to plan their motion on. In this approach, Voxposer [12] focuses on the tabletop manipulation setting and NavCon [3] focuses on creating composable maps for navigation. Using LLM to create reward functions is also promising. Eureka [23, 24] and Language to Rewards for Robotic Skill Synthesis [41] both show that LLM can generate good reward functions that allows robots to acquire complex skills. Finally, LLM can also be used to generate programs for high-level planning. LLM+p [18] outputs a robot plan in the form of the well-defined planning domain definition language (PDDL). Tidybot [39] uses an LLM to generate a rule that captures user preferences from examples and executes a program to sequentially complete the task in order. RoboEval [10] focuses on generating domain-specific programs for service mobile robots. It generates a program that allows the service robot to carry out long-horizon tasks and then validates the correctness of the program.
4.2 Generating Datasets For Fine-tuning LLMs
To enhance LLMs’ performance in code generation, numerous studies have explored the creation of specialized datasets [26, 13, 25]. Self-Instruct [36] is one popular method for generating synthetic datasets using an LLM. Following this methodology, Alpaca [32] generates 52K instruction-following demonstrations and subsequently fine-tunes the LLaMA 7B model [34] to create Alpaca 7B, which can behave qualitatively similarly to OpenAI’s text-davinci-003. Code Alpaca [5] extends this approach to generate code instructions using 21 seed tasks, while Gorilla-LM [29] adapts the method to focus on ML domain-specific APIs from Huggingface, TensorFlow Hub, and Torch Hub. To create more complex instructions, Evol-Instruct [40, 22] proposes iteratively updating instructions to become more complex through different prompting strategies. In addition to Evol-Instruct, OSS-Instruct [38] uses open-source code snippets to generate 75K high-quality instruction data and fine-tunes the Codelllama-Python-7B model to create Magicoder, which can match the performance of GPT-3.5-Turbo [27] on HumanEval [6]. While these works focus on creating seed instruction sets to generate synthetic data for effectively fine-tuning an LLM, our research investigates post-processing methods in addition to Self-Instruct. Specifically, we concentrate on generating domain-specific programs in robotics [10], where we can effectively leverage constraints to filter out erroneous programs.
5 Conclusion, Limitation and Future Works
In this work, we introduce Robo-Instruct, a novel framework to generate synthetic training data to fine-tune small language models for domain-specific robot programs. Robo-Instruct comprises two novel components: 1) RoboSim, an angelic-execution-based algorithm to effectively validate Self-Instruct-generated programs, and 2) InstAlign, an instruction alignment procedure to revise instructions to better align with the generated programs. The experimental results demonstrate that the Codellama-Python-7B model fine-tuned on the Robo-Instruct-generated dataset can significantly outperform many popular open-weight LLMs for generating domain-specific robot programs. It also outperforms two proprietary LLMs, GPT-3.5-Turbo and Gemino-1.0-Pro, as well as the Self-Instruct-fine-tuned variant. A limitation of this study is that Robo-Instruct relies on Self-Instruct to filter invalid programs, making the dataset quality dependent on Self-Instruct’s performance. This can introduce biases if Self-Instruct consistently fails in certain areas. Future work will explore integrating Robo-Instruct with advanced methods like Evol-Inst and OSS-Inst to enhance dataset quality for domain-specific robot programs.
Appendix A Appendix
A.1 Overview
In this appendix, we first present ablation experiments to investigate the percentage of invalid programs generated by Self-Instruct and examine how the generation temperature in InstAlign affects final performance. Next, we analyze and compare the datasets generated by Robo-Instruct and Self-Instruct. Finally, we list the seed tasks used in RoboEvaland the CoT prompt.
A.2 Ablation Exmperiments
Figure 5: Ablation Experiments
A.2.1 effectivenss of the simulator
We analyze the percentage of instruction-program pairs discarded by RoboSim at various maximum resampling limits, as shown in Fig. 5. Initially, with the maximum resampling limit set to 0, disabling the rejection-sampling method, approximately 51% of the programs generated by Self-Instruct contain errors. As the limit increases, fewer programs are discarded. However, there is a diminishing return; even with the maximum resampling limit set to 10, about 15% of the instructions still result in invalid programs.
A.2.2 Instruction Alignment model temperature
We further investigate how varying LLM temperatures for generating $\mathcal{I}_{\text{revised}}$ in InstAlign impact the performance of the fine-tuned model. Fig. 5 shows the bar chart of the pass@1 score of the models fine-tuned over datasets generated using different LLM temperatures in InstAlign. The model performs the best when fine-tuned on the dataset generated using LLM temperature $T=0.3$. As the temperature increases, we observe a decrease in performance.
A.3 Analysis of the Generated Datasets
(a) Token Length Distribution for Self-Instruct vs.Robo-Instruct
(b) Cosine Similarity with RoboEvalfor Self-Instruct vs.Robo-Instruct
Figure 6: Dataset Analysis
Method Size Ngram=4 Score # Synth. Loc. # Synth. Obj.
Robo-Instruct 5K 0.587 1025 928
Self-Instruct 5K 0.581 956 1060
Table 3: Dataset Statistics
We first compute and plot the distribution of token lengths in the Self-Instruct-generated dataset and the Robo-Instruct-generated dataset, as shown in Fig. 6(a). Next, we measure the cosine similarity between each dataset and the RoboEvalbenchmark tasks following the approach in Magicoder [38], as depicted in Fig. 6(b). Finally, Tab. 3 presents the n-gram diversity score of each dataset, along with the number of synthesized locations and objects. Our findings indicate that both distributions and dataset statistics are very similar, suggesting that Robo-Instruct enhances the quality of the generated data over Self-Instruct rather than merely aligning the dataset towards the benchmark tasks.
A.4 RoboEvalSeed Task Example
1# Instruction: Go to Arjun’s office,
2# ask him if he is ready to head out,
3# and come back and tell me what he said
4def task_program():
5 start_loc = get_current_location()
6 go_to("Arjun’s office")
7 response = ask("Arjun",
8 "Are you ready to go?",
9 ["Yes", "No"])
10 go_to(start_loc)
11 say("Arjun said: " + response)
Listing 1: Seed Task Example 1
1# Instruction: Ask Alice if she needs 1, 2, or 3 boxes.
2# Go to the storage room and ask if they have that many boxes.
3# If so, go place the boxes in Alice’s office.
4# Otherwise, tell Alice you could not get the boxes.
5def task_program():
6 go_to("Alice’s office")
7 num_boxes = ask("Alice",
8 "How many boxes do you need?",
9 ["1", "2", "3"])
10 go_to("storage room")
11 response = ask("",
12 "Do you have" + num_boxes + " boxes?",
13 ["Yes", "No"])
14 if response == "Yes":
15 for _ in range(int(num_boxes)):
16 pick("box")
17 go_to("Alice’s office")
18 place("box")
19 go_to("storage room")
20 else:
21 go_to("Alice’s office")
22 say("I could not get the boxes")
Listing 2: Seed Task Example 2
1# Instruction: Check if there is a red marker in the main
2# office, and if so, tell Eve that there is a marker there.
3# If not, go to the supply room and
4# bring a red marker to the main office.
5def task_program():
6 go_to("main office")
7 red_marker_found = is_in_room("red marker")
8 if red_marker_found:
9 go_to("Eve’s office")
10 say("There is a red marker in the main office")
11 else:
12 go_to("supply room")
13 pick("red marker")
14 go_to("main office")
15 place("red marker")
Listing 3: Seed Task Example 3
1# Instruction: Check every classroom if there is a whiteboard.
2# Go to Aiden’s office to tell him which room does not
3# have a whiteboard. Come back and tell me task is completed.
4def task_program():
5 start_loc = get_current_location()
6 list_of_rooms = get_all_rooms()
7 room_without_whiteboard = []
8 for room in list_of_rooms:
9 if "classroom" not in room:
10 continue
11 go_to(room)
12 if not is_in_room("whiteboard"):
13 room_without_whiteboard.append(room)
14 go_to("Aiden’s office")
15 if len(room_without_whiteboard) > 0:
16 message = ""
17 for room in room_without_whiteboard:
18 message += room + ", "
19 message += "do not have a whiteboard"
20 else:
21 message = "all classrooms have a whiteboard"
22 say(message)
23 go_to(start_loc)
24 say("task is completed")
Listing 4: Seed Task Example 4
1# Instruction: Go to the kitchen and wait for someone
2# to show up. When someone shows up, ask them to open
3# the fridge, then pick up a diet coke.
4# Finally, put the diet coke in the living room.
5def task_program():
6 go_to("kitchen")
7 while True:
8 if is_in_room("person"):
9 response = ask("",
10 "Please open the fridge",
11 ["Yes", "No"])
12 if response == "Yes":
13 pick("diet coke")
14 break
15 time.sleep(1)
16 go_to("living room")
17 place("diet coke")
Listing 5: Seed Task Example 5
1# Instruction: Take a bed sheet from the laundry room
2# and put it in each of the bedrooms.
3def task_program():
4 start_loc = get_current_location()
5 list_of_rooms = get_all_rooms()
6 for room in list_of_rooms:
7 if "bedroom" not in room:
8 continue
9 go_to("laundry room")
10 pick("bed sheet")
11 go_to(room)
12 place("bed sheet")
13 go_to(start_loc)
Listing 6: Seed Task Example 6
A.5 Prompts to Generate Synthetic Dataset Using Self-Instruct
You are a helpful assistant. Here is a robot that has the following capabilities:
- def get_current_location() -> str:
- def get_all_rooms() -> list[str]:
- def is_in_room(object : str) -> bool:
- def go_to(location : str) -> None:
- def ask(person : str, question : str, options: list[str]) -> str:
- def say(message : str) -> None:
- def pick(obj: str) -> None:
- def place(obj: str) -> None:
Generate an interesting robot task that can be accomplished using the above capabilities.
{{SEED EXAMPLE}}
Generate an interesting robot task that can be accomplished using the above capabilities.
Table 4: Prompts to Generate Synthetic Dataset Using Self-Instruct.
A.6 CoT Prompts for InstAlign
### Role: You are an expert at understanding robot programs. You will be given a task instruction and robot program pair. However, the instruction may not align with the program well. You need to correct the task instruction to match the given robot program.
### Context: The robot only has access to the following 8 APIs and standard Python functions
- def get_current_location() -> str:
- def get_all_rooms() -> list[str]:
- def is_in_room(object : str) -> bool:
- def go_to(location : str) -> None:
- ask(person : str, question : str, options: list[str]) -> str:
- say(message : str) -> None:
- def pick(obj: str) -> None:
- def place(obj: str) -> None:
### Inputs
Original Instruction: This is a task instruction that may not align with the robot program Robot Program: This is a python function starting with ‘def task_program():‘
### Task:
1. Write down all the provided APIs used in the program and explain the effect of each API in this program
2. Examine these APIs and write down step by step what the program does
3. Combine all the results above and rewrite the instruction under # Final Corrected Instruction: You need to be specific and clear in your final corrected instruction.
Table 5: CoT Prompts for InstAlign.

### v2

@@
 Abstract
-Large language models (LLMs) have shown great promise at generating robot programs from natural language given domain-specific robot application programming interfaces (APIs).
-However, the performance gap between proprietary LLMs and smaller open-weight LLMs remains wide.
-This raises a question: Can we fine-tune smaller open-weight LLMs for generating domain-specific robot programs to close the performance gap with proprietary LLMs?
-While Self-Instruct is a promising solution by generating a diverse set of training data, it cannot verify the correctness of these programs.
-In contrast, a robot simulator with a well-defined world can identify execution errors but limits the diversity of programs that it can verify.
-In this work, we introduce Robo-Instruct, which brings the best of both worlds — it promotes the diversity of Self-Instruct, while providing correctness of simulator-based checking.
-Robo-Instruct introduces RoboSim to synthesize a consistent world state on the fly by inferring properties relevant to the program being checked, and simulating actions accordingly.
-Furthermore, the instructions and programs generated by Self-Instruct may be subtly inconsistent — such as the program missing a step implied by the instruction.
-Robo-Instruct further addresses this with InstAlign, an instruction-program alignment procedure that revises the task instruction to reflect actual results of the generated program.
-Given a few seed task descriptions and the robot APIs, Robo-Instruct is capable of generating a training dataset using only a small open-weight model.
-This dataset is then used to fine-tune small open-weight language models, enabling them to even exceed the performance of several proprietary LLMs including GPT-3.5-Turbo and Gemini-Pro.
-Figure 1: High-Level Overview of Robo-Instruct.
-This figure also illustrates an example of an invalid Self-Instruct-generated instruction and program, as well as pass@1 results of different LLMs on RoboEval.
+Open-weight LLMs are particularly appealing choices to generate training data for fine-tuning Code LLMs on domain-specific service robot applications because they are cost-effective, customizable, and offer better privacy protection.
+However, unlike proprietary LLMs, open-weight models are more error-prone and often produce programs that violate domain-specific constraints.
+A promising solution is to incorporate a robot simulator with a well-defined environment to verify program correctness.
+Yet, these environments require pre-enumeration of relevant entities and their states, which limits the diversity of programs that can be effectively verified.
+In this work, we introduce Robo-Instruct that preserves the diversity of programs generated by an LLM while providing the correctness of simulator-based checking.
+Robo-Instruct introduces RoboSim to dynamically synthesize consistent simulation environments for each generated program.
+Moreover, Robo-Instruct handles subtler instruction-program inconsistencies that do not result in a constraint violation via InstAlign, an LLM-aided instruction-program alignment process.
+Given domain-specific APIs and a few seed examples, Robo-Instruct can leverage an 8B Llama3 model to generate a training dataset for fine-tuning a 7B CodeLlama model.
+Our fine-tuned model achieves a 28.75% improvement in pass@1 over the original base model and a 13.75% improvement compared to its Self-Instruct-finetuned counterparts, even surpassing the performance of a few proprietary LLMs, such as GPT-3.5-Turbo and Gemini-Pro.
+Figure 1: High-level overview of the Robo-Instruct framwork. This figure also shows the pass@1 score performance of Robo-Instruct fine-tuned LLM compared to other LLMs on RoboEval.
 1 Introduction
-Large language models (LLMs) have demonstrated great promise at generating robot programs from natural language instructions [10, 12, 3, 18, 39, 17, 31, 11].
-For example, consider an instruction for a service mobile robot: "Check how many conference rooms have no markers." The robot may be equipped with a domain-specific robot application programming interface (API) that includes skills such as go_to(location) for navigation and is_in_room(object) for perception.
-Since such domain-specific APIs do not exist in the training dataset of general-purpose LLMs, in-context learning (ICL) via few-shot examples is often employed to describe and use such APIs for performing few-shot inference.
-However, there is a significant performance gap [10] in the correctness of programs generated by ICL for large proprietary models and smaller open-weight models that can be deployed locally on robots.
-This raises a question: can we fine-tune small open-weight LLMs for generating domain-specific robot programs to close the performance gap with proprietary LLMs?
-Since training datasets of the domain-specific robot programs are often unavailable, Self-Instruct might seem like a promising solution [36, 29].
-Consider the setting of generating programs for service mobile robots that can perceive objects, navigate to various locations, manipulate items, and communicate with humans.
-By formulating these robot skills into APIs, we can create a few seed task examples demonstrating their use case and employ Self-Instruct to generate a diverse set of instruction-program pairs as training data, as illustrated in Fig.
-1.
-However, using Self-Instruct naïvely may generate infeasible instructions—e.g., asking the robot to pick up multiple objects at once when it cannot due to physical constraints.
-They can also violate domain-specific constraints.
-For example, in Fig.
-1, after line 2 confirms the absence of a key at the current location, line 3 erroneously attempts to pick up a key.
-Further, these instructions may not align with the generated programs, even if these programs are valid.
-For example, Fig.
-1 shows an example instruction directing the robot to verbally ask in each room if a key exists, whereas the program instructs the robot to visually check in each room.
-Finally, the generated programs may have execution errors.
-These challenges may appear to be solvable using a simulator, but a simulator needs an initial world state to check against programs.
-A simulator using a hand-curated world state will end up rejecting the wide diversity of programs generated by Self-Instruct, even if they are executable, just because the world state did not capture some aspect relevant to them (e.g., the presence of a “key”).
-This work introduces Robo-Instruct, a new framework based on Self-Instruct, to address these issues and improve the performance of small open-weight language models for generating domain-specific robot programs.
+Large language models (LLMs) have shown great promise at leveraging domain-specific application programming interfaces (APIs) and generating robot programs from natural language instructions (Hu et al., 2024; Huang et al., 2023b; Biggie et al., 2023; Liu et al., 2023a; Wu et al., 2023; Liang et al., 2022; Singh et al., 2023; Huang et al., 2023a).
+For instance, by formulating a robot’s navigation and perception skills into APIs, such as go_to(location) and is_in_room(object), an LLM can generate a program for a service mobile robot to complete the task: “Determine the number of conference rooms without markers”.
+However, despite impressive results, the performance gap remains wide between proprietary and open-weight LLMs in generating robot programs from domain-specific APIs (Hu et al., 2024).
+To bridge the performance gap, Self-Instruct (Wang et al., 2022) is a popular method for generating domain-specific data for finetuning LLMs.
+Given the definition of APIs and a few seed task examples, Self-Instruct prompts an LLM to generate diverse instruction-program pairs as training data.
+Figure 2: Examples of programs violating domain-specific constraints.
+This makes Self-Instruct a seemingly appealing way to prompt open-weight LLMs to generate training data, for the fine-tuning of Code LLMs on domain-specific service robot applications (due to their cost-effectiveness, better privacy protection, and customizability).
+However, open-weight LLMs are still prone to errors.
+As a result, using Self-Instruct naively can produce low-quality data, such as programs that violate domain-specific constraints or instructions that are infeasible for the robot to execute.
+For instance, as shown in the figure Fig.
+2, the program may attempt to pick up an apple that is not present in the current location (example 1); or the instruction might ask the robot to pick up multiple objects simultaneously, which is physically impossible (example 2).
+To catch these errors, executing the program in a simulator with well-defined environments may appear promising.
+However, such environments require pre-enumeration of relevant entities and their states, which can result in rejecting a vast number of programs generated by Self-Instruct even if they are executable, simply because the pre-defined simulation environment fails to capture certain relevant aspects (e.g., the presence of an “apple”); this undermines the diversity of the data generated.
+This work introduces Robo-Instruct to bring the best of both worlds — it preserves the diversity of programs generated by an LLM while providing the correctness of simulator-based checking.
+Robo-Instruct is a new framework based on Self-Instruct, to address the challenges posed by using an open-weight LLM for generating domain-specific robot programs.
 As shown in Fig.
 1, Robo-Instruct introduces two novel components: (1) RoboSim, a task-agnostic simulator that encodes domain-specific constraints and validates robot programs generated from Self-Instruct.
-Critically, RoboSim dynamically synthesizes a consistent world state starting from arbitrary programs.
-(2) InstAlign, an instruction-program alignment procedure that revises the generated instructions to better reflect the intent of the generated programs.
-Robo-Instruct also employs a rejection-sampling mechanism that rejects invalid programs detected by RoboSim and queries Self-Instruct for a new program corresponding to the same generated instruction.
-We validate Robo-Instruct by fine-tuning Codellama-Python-7B [30] and evaluate on RoboEval, a domain-specific code generation benchmark for service mobile robots.
-We show that Robo-Instruct is capable of improving the performance of the Codellama model by using only a small open-weight model to generate the training dataset.
-Compared to the base Codellama-Python-7B model without fine-tuning, our Robo-Instruct fine-tuned models outperform by $28.75\%$ in average pass@1 scores; and, compared to Self-Instruct fine-tuned model, our model outperform by $13.75\%$.; and the best pass@1 of Robo-Instruct fine-tuned model achieves a 68.75% match, surpassing the performance of the proprietary GPT-3.5-Turbo and Gemini-1.0-Pro.
+Critically, RoboSim dynamically synthesizes consistent simulation environments starting from arbitrary programs.
+(2) InstAlign, an LLM-aided instruction-program alignment procedure that revises the generated instructions to better reflect the intent of the generated programs.
+Robo-Instruct also uses a rejection-sampling mechanism to discard invalid programs detected by RoboSim and query Self-Instruct for a new program based on the same instruction.
+We use an 8B Llama3 model (AI, 2024) to generate instruction-program training data and fine-tune a 7B CodeLlama model (Rozière et al., 2024) on RoboEval, a domain-specific benchmark for service robot programs.
+Our Robo-Instruct-fine-tuned model significantly outperforms the base model by 28.75% in pass@1 scores and achieves a 13.75% improvement over its Self-Instruct-fine-tuned variant.
+Moreover, it surpasses several larger code models, including Deepseek-Coder-33B (Guo et al., 2024), Starcoder2-15B (Lozhkov et al., 2024), GPT-3.5-Turbo (OpenAI, 2022), and Gemini-Pro (Team et al., 2024).
+Finally, in real-world deployment, we demonstrate that Robo-Instruct achieves significantly lower latency compared with GPT-3.5 and GPT-4.
 Contributions
-Our main contributions are as follows:  [^1] Website coming soon: https://amrl.cs.utexas.edu/robo-instruct/
+Our main contributions are as follows  [^1] Project Website: https://amrl.cs.utexas.edu/robo-instruct/
   1. 1.
-    We introduce Robo-Instruct, a new framework for improving the code generation performance of small open-weight language models for domain-specific robot programs.
+    We introduce Robo-Instruct, a new framework for improving open-weight LLMs’ ability to generate training data for finetuning Code LLMs on domain-specific service robot applications.
 This framework introduces two novel components, RoboSim and InstAlign.
   2. 2.
-    We introduce a dynamic world synthesis and evaluation process for generating relevant world states for automated code checking for diverse, arbitrary tasks in RoboSim.
+    We present RoboSim, a method to synthesize consistent simulation environments dynamically for verifying the generated programs.
+This method preserves the diversity of programs generated by an LLM while providing correctness with simulator-based checking.
   3. 3.
-    We introduce InstAlign, an instruction alignment procedure to refine instruction-code pairs to improve alignment between instructions and code generated by Self-Instruct.
+    We present InstAlign, a simple instruction-program alignment procedure that revises generated instructions better to reflect the actual results of the generated program.
   4. 4.
-    We fine-tune a small open-weight model, Codellama-Python-7B [30], using Robo-Instruct, and improve its performance to outperform several CodeLLMs, including Deepseek-Coder-33B [8], and Starcoder2-15B [21] and two proprietary LLMs, GPT-3.5-Turbo [27] and Gemini-1.0-Pro [33] on the RoboEval benchmark.
+    We show that the Robo-Instruct finetuned model can significantly outperform the original base model and its finetuned variant using Self-Instruct in generating domain-specific service robot programs.
+It also surpasses several other larger Code LLMs.
+  5. 5.
+    We demonstrate that the Robo-Instruct fine-tuned model has significantly faster inference than proprietary models, making it more suitable for real-world robot deployments.
 2 Robo-Instruct
-In this section, we present how Robo-Instruct generates training datasets of domain-specific robot programs.
-Alg.
-1 shows a broad overview of the framework.
-To add an entry in the training dataset, Self-Instruct first generates an instruction-program pair, $(\mathcal{I},\mathcal{P})$, from the robot APIs and seed tasks, shown in Appendix A.4.
-Then, RoboSim dynamically synthesizes a consistent world state on the fly as it executes and validates $\mathcal{P}$.
-If $\mathcal{P}$ is invalid, Robo-Instruct employs a rejection-sampling method, which generates a new program $\mathcal{P}$ given the same $\mathcal{I}$ and evaluates the new $\mathcal{P}$ again.
-This process repeats until $\mathcal{P}$ becomes valid or a predefined maximum resampling limit is reached.
-If the limit is reached, the instruction might be invalid given the domain-specific APIs or too complex to generate a program, so the instruction-program pair is discarded.
-Finally, if $\mathcal{P}$ is valid, InstAlign takes in $(\mathcal{I},\mathcal{P})$ to revise $\mathcal{I}$ to better reflect the intent of $\mathcal{P}$ and the aligned instruction and program is saved to the training dataset.
-In the following subsections, we elaborate on the specific design of each component.
-Algorithm 1 Robo-Instruct: Instruction-Program Generation
+In this section, we describe how Robo-Instruct generates training datasets for domain-specific service robot programs.
+Fig.
+1 provides a high-level overview of the framework.
+To add an entry to the training dataset, Self-Instruct first generates an instruction-program pair based on robot APIs and seed tasks (detailed prompts in Appendix A.4.2).
+Then, RoboSim dynamically synthesizes a consistent simulation environment to validate the program as it executes.
+If the program is invalid, Robo-Instruct applies a rejection-sampling method, which generates a new program based on the same instruction and re-evaluates it.
+This process continues until the program is valid or a predefined maximum resampling limit is reached.
+If the limit is exceeded, the instruction may be incompatible with the domain-specific APIs or too complex, so the instruction-program pair is discarded.
+Finally, if the program is valid, InstAlign uses an open-weight LLM to refine the instruction to better align with the program’s intent, and the revised instruction-program pair is added to the training dataset.
+In the following subsections, we delve into the specific design of each component.
+2.1 RoboSim: Dynamic Synthesis of Simulation Environments for Program Validation
+We present a principled approach to designing RoboSim, a system for dynamically synthesizing consistent simulation environments to validate domain-specific robot programs.
+For service mobile robots, a simulation environment often relies on three concepts:
+  1. 1.
+    A list of entities to reason about, e.g., “apple”, “kitchen”
+  2. 2.
+    The type of the entities, and hence their affordances, e.g., “apple” is an object, you can pick it up; “kitchen” is a location, you can go to it, and it contains objects.
+  3. 3.
+    The state of the entities in the world, e.g., the “apple” is in the “kitchen”.
+These concepts are closely related to the domain-specific APIs, where each API invocation during program execution can trigger updates to the simulation environment.
+To handle the interaction between APIs and simulation environments, we introduce DynamicEval, an algorithm inspired by Angelic Execution (Broy and Wirsing, 1981), a software engineering technique to infer program properties from incomplete API specifications.
+DynamicEval automatically generates a simulation environment for each program and checks its correctness within the inferred environment.
 […]
-1:  Initialize: $\mathcal{D}=\emptyset$triangleright Training dataset
-2:  Initialize: $N$triangleright Training dataset size
-3:  Initialize: $m$triangleright Maximum resampling limit
-4:  while $\text{len}(\mathcal{D})<N$ do
-5:     $\mathcal{I}\leftarrow\textsc{Self-Instruct}_{inst}(\mathcal{S})$
-6:     $\mathcal{P}\leftarrow\textsc{Self-Instruct}_{code}(\mathcal{S},\mathcal{I})$
-7:     for $i=1$ to $m$ do
+1:  Input: api_fn $\triangleright$ The API function name
+2:  Input: api_inputs $\triangleright$ The input received by the API invocation
+3:  Input: $\mathcal{W}$triangleright The current simulation environment
+4:  info $\leftarrow$ Extract_API_Invocation_Info(api_fn, api_inputs)
+5:  entities $\leftarrow$ Infer_Relevant_Entities(info)
+6:  for entity $\in$ entities do
+7:     if Is_Entity_Initialized(entity, $\mathcal{W}$then
+8:        entity_type $\leftarrow$ Deduce_Type(info)
+9:        entity_new_state $\leftarrow$ Deduce_State(info)
 […]
-[…]
-10:           $\mathcal{P}\leftarrow\textsc{Self-Instruct}_{code}(\mathcal{S},\mathcal{I})$triangleright Rejection-sampling
-11:        else
-12:           $\mathcal{I}_{\text{aligned}}\leftarrow\textsc{InstAlign}(\mathcal{S},\mathcal{I},\mathcal{P})$triangleright Align instruction with the program
-13:           $\mathcal{D}\leftarrow(\mathcal{I}_{\text{aligned}},\mathcal{P})$
-14:           break
-15:        end if
-16:     end for
-17:  end while
-18:  return  $\mathcal{D}$
-2.1 RoboSim: A Task-Agnostic Simulator For Domain-Specific Programs
+11:           Update_Entity_State(entity, info, $\mathcal{W}$)
+12:        else
+13:           raise “Error: state inconsistent or type mismatch”
+14:        end if
+15:     else
+16:        Initialize_Entity_With_Random_State(entity, info, $\mathcal{W}$)
+17:     end if
+18:  end for
+19:  retval $\leftarrow$ Get_Return_Value(info, $\mathcal{W}$)
+20:  return  retval
+As shown in Alg.
+1, upon each API invocation, the corresponding inputs and the current simulation environment are passed into DynamicEval.
+DynamicEval first infers relevant entities when they appear in the program being checked (line 5).
+For instance, if a program includes the statement pick(”apple”), DynamicEval infers that apple is an entity to consider, even if it is not currently defined in the simulation environment.
+For each relevant entity, if it has already been initialized (line 7), DynamicEval infers its type and new state from the API invocation and checks for any inconsistencies with the current simulation environment (lines 8-10).
+For example, pick requires an object type, while go_to requires a location type  [^2] In this example, type compatibility check is strict (i.e., “apple” is only an object and no further inference is made about its location).
+Nevertheless, the algorithm is also capable of handling more advanced scenarios..
+If a program contains:
+1def task_program():
+2 pick("apple")
+3 go_to("apple")
+By invoking API calls sequentially, DynamicEval first infers that “apple” is an object and then raises an error when go_to(“apple”) is called.
+If no inconsistency is detected, the simulation environment is updated accordingly based on the API definition (line 11).
+On the other hand, if the entity has not been initialized, it will be assigned a random plausible state (line 16).
+For example, the API is_in_room(object) checks if an object is in the same location as the robot and expects a boolean return.
+In this case, DynamicEval assigns a 50% probability for the apple to be in the robot’s current location — determining the state of the apple as either present or absent in the robot’s location:
+1def task_program():
+2 if is_in_room("apple"):
+3 ...
+Finally, DynamicEval computes and returns the value based on the API specification and the updated simulation environment.
+Fig.
+3 illustrates how RoboSim leverages DynamicEval to dynamically synthesize the simulation environment and validate the program.
+In RoboSim, the simulation starts with only the robot at its initial position, and entities are added as the program runs.
+At execution timestep 2a, after the robot picks up an apple, it becomes unclear whether another apple remains at the location.
+Therefore, DynamicEval sets the apple’s state to ”Undefined” and removes it in subsequent executions.
+This method is also related to STRIPS planning, as we demonstrate the connection in Appendix A.2.
+Figure 3: Illustration of RoboSim executing a task program and dynamically updating the simulation environment.
+Initially, the environment only contains the robot’s starting location.
+As the program runs, two possible execution paths emerge (purple and blue).
+The environment is updated by DynamicEval at each step, reflecting the robot’s actions, object states, and location changes along both paths.
+In the example program shown in Fig.
+3, it’s clear to humans that the program’s logic is flawed because it attempts to pick up the apple, which is not present in the room.
+But how would the simulator identify this as a failing state?
+RoboSim solves this issue by simulating all possible states of the discovered entities and verifying that none lead to erroneous program execution.
+In this case, the “apple” can either be present in the room or not.
+If the apple is not present, executing pick(”apple”) will result in an error.
+Checking all possible states requires exploring an exponentially growing number of combinations based on the entities discovered.
+To manage this, RoboSim employs a bounded compute budget to randomly sample from this exponential space, as detailed in the overall RoboSim algorithm in Alg.
+2.
 Algorithm 2 $\textsc{RoboSim}(\mathcal{P})$
@@
 2:  Initialize: $k$triangleright Number of evaluation iterations
-3:  Initialize: $\mathcal{W}_{\text{init}}$triangleright An initial world state with or without predefined information
+3:  Initialize: $\mathcal{W}_{\text{init}}$triangleright An initial simulation environment with or without predefined information
 4:  $\mathcal{P}_{\text{trans}}\leftarrow\textsc{Translate}(\mathcal{P},\mathcal{A},\textsc{DynamicEval})$triangleright Replace each API call with DynamicEval
@@
 6:   try:
-7:    $\mathcal{W}\leftarrow\mathcal{W}_{\text{init}}$triangleright Initialize a new world state
+7:    $\mathcal{W}\leftarrow\mathcal{W}_{\text{init}}$triangleright Initialize a new simulation environment
 8:    exec$(\mathcal{P}_{\text{trans}},\mathcal{W})$
@@
 12:  return  True $\triangleright$ Return True if all program executions are successful
-[…]
-[…]
-2:  for $l\in p$ do $\triangleright$ Loop through every literal in the precondition
-3:   if $\textsc{CheckDefined}(\mathcal{W},l)==\text{undefined}$ then
-4:    $W\leftarrow\textsc{GrowWorld}(l,\mathcal{W})$triangleright Instantiate the literal and grow $\mathcal{W}$ to include it
-5:   end if
-6:  end for
-7:  $\text{retval},\mathcal{W}\leftarrow\textsc{ExecUpdate}(\text{api\_fn},\text{params},\mathcal{W})$triangleright Execute api_fn and update $\mathcal{W}$
-8:  return  retval, $\mathcal{W}$
-We present a principled approach to design RoboSim for validating domain-specific robot programs.
-Alg.
-2 illustrates the high-level algorithm used to assess the correctness of a robot program.
-RoboSim employs the concept of world state to simulate the robot actions directed by a program, ensuring consistent and reliable evaluation.
-A world state is a symbolic representation of the environment in which the robot operates, and it keeps track of the high-level changes in the robot state and the surrounding environment as the robot performs actions in order.
-For example, consider a program instruction that commands a robot to check if a pen is nearby.
-The world state queries the stored information about the surrounding environment, identifies all objects at the robot’s current location, and informs the program whether a pen is present.
-However, since Self-Instruct generates arbitrary programs based on the provided APIs, RoboSim does not know what a plausible world state relevant to the program would be a priori — e.g., reasoning about the existence of a pen in the example program shown in Fig.
-1.
-Thus, we equip RoboSim with the ability to synthesize the world state as more robot actions are performed.
-Our approach is inspired by angelic execution [4], which has previously been used for software verification of programs with partially defined library functions.
-In our case, instead of partially defined library functions, we have unknown plausible world states.
-RoboSim dynamically synthesizes and grows a world state based on domain-specific constraints (e.g., object permanence, robot skills, etc.) and the execution trace of the program, which allows it to infer a consistent and relevant world state.
-Specifically, RoboSim modifies the program to replace all API calls with the DynamicEval function (Alg.
-2 line 4) — when an API function is called during execution, the DynamicEval function is invoked instead.
-DynamicEval makes an important extension to the formulation of STRIPS [7] to integrate with API functions.
-DynamicEval equips each API function with specific pre-conditions, effects, and return values.
-The pre-conditions are composed of literals tailored to the function’s requirements.
-For instance, the API function is_in_room("pen"), which determines if an object ‘pen’ is in the same room as the robot, uses two literals for its pre-condition: robot_at(X) and obj_at(X, "pen").
-Generally, STRIPS assigns one of two possible values to each literal: True if the literal is defined, otherwise False.
-However, prior to program execution, DynamicEval is unaware of the program-relevant literals.
-Thus we assign a third value, undefined, to such unknown literals.
-Literals must thus be explicitly defined as either True or False, or they remain undefined if not specified.
-Alg.
-3 demonstrates how DynamicEval executes an API function and updates the world state.
-First, it calculates the precondition specified for the function.
-It then checks each literal in the precondition to see if it is defined.
-If a literal is undefined, DynamicEval invokes GrowWorld, a stochastic function that assigns a random truth value to the literal and updates the world state accordingly.
-Finally, DynamicEval proceeds to execute the API function using the current world state, retrieves the return values, and applies the function’s effects to update the world state.
-Fig.
-2 illustrates an example of RoboSim executing a generated program.
-Initially, RoboSim’s world state only specifies the robot’s current location, and whether a pen is in the same room as the robot remains undefined (line 2).
-Therefore, DynamicEval invokes GrowWorld to randomly determine a truth value for the obj_at(start_loc, "pen") literal, leading to two distinct execution paths depicted in light purple and blue.
-Subsequently, as additional API functions are called, more literals are introduced or updated in the world state to ensure consistent evaluations.
-In addition, a crucial aspect of RoboSim’s design is that a literal can also be set to an undefined value.
-In world state 2 on the purple execution path, after executing the pick("pen") function, the program is unable to determine if another pen exists in the current room, given the provided APIs.
-As a result, the world state updates the literal obj_at(start_loc, "pen") from True to Undefined.
-Finally, due to the stochastic nature of DynamicEval, RoboSim must execute the generated program multiple times to validate the program.
-If all executions are successful, the program is deemed correct (Alg.
-2 line 5-11).
-Figure 2: Example of RoboSim executing a generated program and updating the world state.
-Initially, RoboSim begins with a world state that includes only the robot’s current location.
-As the program executes, two distinct execution paths emerge, depicted in light purple and blue.
-This figure demonstrates how the world state is updated along each execution path.
-2.2 InstAlign: Instruction-Program Alignment Procedure
-Given that LLMs are extensively trained in code understanding [30], InstAlign is a procedure that prompts an LLM to revise $\mathcal{I}$ to better reflect the intent of $\mathcal{P}$.
-This procedure involves two steps: first, given $\mathcal{I}$ and $\mathcal{P}$, InstAlign leverages Chain-of-Thought reasoning [37] (CoT) to prompt an LLM to generate a revised instruction, $\mathcal{I}_{\text{revised}}$; second, InstAlign invokes the LLM again to determine whether $\mathcal{I}$ or $\mathcal{I}_{\text{revised}}$ is more aligned with $\mathcal{P}$’s intent and output the chosen instruction as $\mathcal{I}_{\text{aligned}}$.
-To generate $\mathcal{I}_{\text{revised}}$, the prompt to the LLM comprises the robot API function definitions, $\mathcal{I}$, $\mathcal{P}$, and CoT instructions.
-The CoT asks the LLM to perform the following three steps in order: 1.
-write down all the robot APIs used in the program; 2.
-examine these APIs and write down step by step what the program does; 3.
-combine all the information above to revise the robot instruction.
-Similarly, to determine $\mathcal{I}_{\text{aligned}}$, an LLM is prompted to think step by step about $\mathcal{P}$, $\mathcal{I}$ and $\mathcal{I}_{\text{revised}}$ to arrive at a conclusion.
-Detailed prompt is shown in Appendix A.6.
+2.2 InstAlign: LLM-aided Instruction-Program Alignment Procedure
+Figure 4: Overview of InstAlign.
+Even after RoboSim verifies that a program doesn’t violate domain-specific constraints, it may still have subtle inconsistencies with the instructions — such as omitting a step implied by the instruction.
+The proposed rejection-sampling strategy alone cannot resolve this issue, and correcting verified programs to fully reflect the instructions while ensuring they remain valid is a challenging task.
+We note that LLMs nowadays have demonstrated impressive code understanding capabilities (Rozière et al., 2024; Nam et al., 2024; Leinonen et al., 2023; Li et al., 2023; Lekshmi-Narayanan et al., 2024).
+Instead of correcting the program, it may be more effective for the LLM to revise its generated instructions to better align with the program’s intent.
+We propose InstAlign, a procedure that prompts an LLM to revise generated instructions to better match the intent of the program, as shown in Fig.
+4.
+InstAlign follows two steps: first, it uses Chain-of-Thought reasoning (CoT) (Wei et al., 2022) to generate a revised instruction based on the instruction-program pair; then, it prompts the LLM to compare the original and revised instructions, selecting the one that best aligns with the program’s intent.
+Detailed prompts are provided in Appendix A.4.3.
 3 Analysis and Experiments
@@
   1. 1.
-    Is Robo-Instruct effective at generating training data to fine-tune a small language model for generating domain-specific robot programs?
+    Is Robo-Instruct effective at generating training data to fine-tune a small language model to generate domain-specific programs for robots?
   2. 2.
     How do RoboSim and InstAlign impact the effectiveness of Robo-Instruct?
-We conduct our investigation by fine-tuning the Codellama-Python-7B model [30] on the synthetic dataset generated by Robo-Instruct and evaluate the fine-tuned model using RoboEval [10], a domain-specific code generation benchmark for service mobile robots.
+We conduct our investigation by fine-tuning the CodeLlama-Python-7B model (Rozière et al., 2024) on the synthetic dataset generated by Robo-Instruct and evaluate the fine-tuned model using RoboEval (Hu et al., 2024), a domain-specific code generation benchmark for service mobile robots.
 In the following subsections, we first provide a brief description of RoboEval.
@@
 5 illustrates these function signatures and definitions, alongside an example task instruction and its canonical solution from the benchmark.
-In addition, unlike other popular code generation benchmark tasks [6, 2, 16, 19, 14, 9], the order of the robot’s actions is crucial for successfully completing the specified tasks.
-For instance, in the task "bring me a marker from the classroom that does not have a whiteboard," the robot must check each classroom until it finds one without a whiteboard, whereas simply bringing back a marker is insufficient.
+In addition, unlike other popular code generation benchmark tasks (Chen et al., 2021; Austin et al., 2021; Li et al., 2022; Liu et al., 2023b; Lai et al., 2022; Hendrycks et al., 2021), the order of the robot’s actions is crucial for successfully completing the specified tasks.
+For instance, in the task ”bring me a marker from the classroom that does not have a whiteboard,” the robot must check each classroom until it finds one without a whiteboard, whereas simply bringing back a marker is insufficient.
 Hence, RoboEval evaluates the generated program by executing it in a simulator to capture the action traces, which are subsequently validated for sequence correctness using temporal logic.
@@
 - GPT-4 - 83.75% 85.81% Proprietary
-- GPT-3.5 - 67.5% 65.56% Proprietary
-- Gemini-1.0-Pro - 60.00% 59.88% Proprietary
+- GPT-3.5-Turbo - 67.5% 65.56% Proprietary
+- Gemini-Pro - 60.00% 59.88% Proprietary
 - Codellama-Python 7B 40.00% 39.31% Open
@@
 For the LLM used in InstAlign, we empirically adjust the generation temperature to $T=0.3$ to optimize performance.
-Furthermore, we assess the edit similarity between token sequences of each instruction pair in the dataset [15], removing duplicates where the similarity score exceeds 0.6.
+Furthermore, we assess the edit similarity between token sequences of each instruction pair in the dataset (Lee et al., 2022), removing duplicates where the similarity score exceeds 0.6.
 We use the same setup to generate data via Self-Instruct.
@@
 The learning rate is set to be $-53e\text{-}$ 5 with a warmup ratio of $3\%$ and a constant lr scheduler.
-We employ the AdamW optimizer [20] with an effective batch size of 8, training each model for 5 epochs using a sequence length of 2048 tokens.
-We train all our models on a single H-100 GPU using unsloth [35].
+We employ the AdamW optimizer (Loshchilov and Hutter, 2019) with an effective batch size of 8, training each model for 5 epochs using a sequence length of 2048 tokens.
+We train all our models on a single H-100 GPU using unsloth (Unslothai, 2024).
 Baselines.
-We divide our baseline models into 2 categories: 1) proprietary LLMs, including GPT4 [28], GPT3.5-Turbo [27], Gemino-Pro [33], and 2) open-weight LLMs, including Codellama-Python-7B [30], Codellama-Python-34B, Starcoder2-33B [21], Deepseek-Coder-33B [8], and Llama3-8B-Inst [1].
-All the results are evaluated using RoboEval and reported in Tab.
-1.
+We divide our baseline models into 2 categories: 1) proprietary LLMs, including GPT4 (OpenAI et al., 2024), GPT3.5-Turbo (OpenAI, 2022), Gemino-Pro (Team et al., 2024), and 2) open-weight LLMs, including Codellama-Python-7B (Rozière et al., 2024), Codellama-Python-34B, Starcoder2-33B (Lozhkov et al., 2024), Deepseek-Coder-33B (Guo et al., 2024), and Llama3-8B-Inst (AI, 2024).
 Tab.
-1 presents the average pass@1 results for different LLMs on RoboEval, using two different temperature settings for generation: greedy decoding at a temperatures of $T=0$ and nucleus sampling at a temperature of $T=0.2$.
-The results show that Robo-Instruct-fine-tuned Codellama significantly improves upon the base Codellama-Python-7B and outperforms the Self-Instruct-fine-tuned variant.
-Notably, it surpasses all open-weight models, including larger ones like Codellama-Python-34B and Deepseek-Coder-33B.
-Additionally, although the training dataset was generated using Llama3-8B-Inst, which scores less than 50% pass@1 on RoboEval, our Robo-Instruct-fine-tuned model still achieves a significant improvement, scoring 68.75% under deterministic temperature settings for generation.
-Finally, compared to proprietary models, while our Robo-Instruct-fine-tuned model trails the more powerful GPT-4, it outperforms GPT-3.5-Turbo and Gemini-1.0-Pro in generating programs for service mobile robots.
-This result demonstrates the effectiveness of our approach in generating domain-specific robot program data for fine-tuning a small language model.
+1 presents the average pass@1 results for different LLMs on RoboEval, using two settings: greedy decoding at temperature $T=0$ and nucleus sampling at temperature $T=0.2$.
+The results show that Robo-Instruct-fine-tuned CodeLlama significantly improves upon the base CodeLlama-Python-7B, and outperforms the Self-Instruct-fine-tuned variant (Appendix A.3.3 shows that improvements over Self-Instruct are not the result of distributional biases in the selection process).
+Notably, Robo-Instruct surpasses all open-weight models, including larger ones like CodeLlama-Python-34B and Deepseek-Coder-33B.
+Additionally, although the training dataset was generated using Llama3-8B-Inst, which scores less than 50% pass@1 on RoboEval from Tab.
+1, our Robo-Instruct-fine-tuned model still achieves a significant improvement, scoring 68.75%.
+Finally, compared to proprietary models, while our Robo-Instruct-fine-tuned model trails the more powerful GPT-4, it outperforms GPT-3.5-Turbo and Gemini-Pro in generating programs for service mobile robots.
+This result demonstrates the effectiveness of our approach in generating domain-specific robot program data for fine-tuning a much smaller language model.
 It suggests that the fine-tuned model could potentially replace some proprietary models, providing a more cost-effective and private option for local deployment.
@@
 […]
+Figure 6: Self-Instruct-Generated Program Errors.
+Examples highlight errors that violate domain-specific constraints.
+[^4] Programs have been adapted to succinctly demonstrate the types of errors and fit within the figure.
 3.4 Qualitative analysis of the generated program errors
-Figure 4: Self-Instruct-Generated Program Errors: Examples 1 to 4 illustrate errors specific to the Python language, and Examples 5 to 8 highlight errors rooted in domain-specific constraints.
-[^3] Programs have been adapted to succinctly demonstrate the types of errors and fit within the figure.
-We analyze invalid programs identified by RoboSim, categorizing the errors into two types: language-native errors and domain-specific constraint violations.
-Fig.
-3 displays eight examples of these programs, with Examples 1 to 4 illustrating errors specific to the Python language, and Examples 5 to 8 highlighting errors rooted in domain-specific constraints.
-Language-native errors are generally straightforward, such as syntax errors, the use of undefined variables or functions, or improper use of provided APIs.
-In contrast, errors related to domain-specific constraints tend to be more complex to detect.
-For instance, Example 5 illustrates the program incorrectly trying to pick up a watering can (line 3) after establishing that it is not present at the location (line 2).
-Similarly, Example 6 demonstrates an error where the program inappropriately asks Jack (line 5) after confirming his absence from the room (line 3).
-Example 7 illustrates a scenario in which RoboSim updates the world state by labeling "item storage room" as a location after executing the go_to command (line 2).
-Subsequently, the robot attempts to pick up this location (line 3), resulting in an error.
-Example 9 is the most intricate scenario where the world state in the living room is updated to include a toy after the robot places it there (line 7).
-When the robot returns to the living room for the second time (line 5), it does not place down what it holds (line 7).
-Hence, in the third room the robot visits (line 3), when it attempts to pick up a toy again (line 4), an error occurs because the robot can only carry one item at a time.
+We qualitatively analyze invalid programs identified by RoboSim, as shown in Fig.
+4.
+The first three examples are easily recognizable to humans as flawed.
+However, the last example is more complex and involves an error when the robot can navigate to more than two rooms.
+After the robot places a toy in the living room, DynamicEval updates the environment to reflect that a toy is now in the room (line 8).
+However, when the robot returns to the living room later (line 6), it will not drop the item it’s holding (line 8).
+As a result, when the robot enters a third room (line 4) and tries to pick up another toy (line 5), an error will occur because the robot is only capable of carrying one item at a time.
+This example demonstrates that RoboSim can detect invalid programs beyond those easily identifiable through human inspection.
+4 Real-World Deployment Results
+We deployed the Robo-Instruct fine-tuned model to generate and execute mobile robot programs in the real world, as shown in Fig.
+7.
+Compared to GPT-4 and GPT-3.5-turbo (Tab.
+3), our model generates programs about 6x faster than GPT-4 and 3x faster than GPT-3.5-turbo, with similar output quality.
+In Appendix A.5 we showcase more results on long-horizon tasks beyond RoboEval.
+Figure 7: Deployment of the Robo-Instruct fine-tuned model to generate programs based on user-provided instructions and execute them on the robot.
+Models GPT-4 GPT-3.5-Turbo Robo-Instruct (Ours)
+Inference Speed 19 tokens/s 42 tokens/s 114 tokens/s
+Table 3: Inference speed of different models.
 5 Related Work
 5.1 LLMs for Robot Code Generation
-LLMs have shown impressive capabilities in generating robot programs from natural language [17, 31, 11].
+LLMs have shown impressive capabilities in generating robot programs from natural language (Liang et al., 2022; Singh et al., 2023; Huang et al., 2023a).
 One popular approach uses LLMs to generate composable costmaps for robots to plan their motion on.
-In this approach, Voxposer [12] focuses on the tabletop manipulation setting and NavCon [3] focuses on creating composable maps for navigation.
+In this approach, Voxposer (Huang et al., 2023b) focuses on the tabletop manipulation setting and NavCon (Biggie et al., 2023) focuses on creating composable maps for navigation.
 Using LLM to create reward functions is also promising.
-Eureka [23, 24] and Language to Rewards for Robotic Skill Synthesis [41] both show that LLM can generate good reward functions that allows robots to acquire complex skills.
+Eureka (Ma et al., 2023, 2024) and Language to Rewards for Robotic Skill Synthesis (Yu et al., 2023) both show that LLM can generate good reward functions that allows robots to acquire complex skills.
 Finally, LLM can also be used to generate programs for high-level planning.
-LLM+p [18] outputs a robot plan in the form of the well-defined planning domain definition language (PDDL).
-Tidybot [39] uses an LLM to generate a rule that captures user preferences from examples and executes a program to sequentially complete the task in order.
-RoboEval [10] focuses on generating domain-specific programs for service mobile robots.
+LLM+p (Liu et al., 2023a) outputs a robot plan in the form of the well-defined planning domain definition language (PDDL).
+Tidybot (Wu et al., 2023) uses an LLM to generate a rule that captures user preferences from examples and executes a program to sequentially complete the task in order.
+RoboEval (Hu et al., 2024) focuses on generating domain-specific programs for service mobile robots.
 It generates a program that allows the service robot to carry out long-horizon tasks and then validates the correctness of the program.
 5.2 Generating Datasets For Fine-tuning LLMs
-To enhance LLMs’ performance in code generation, numerous studies have explored the creation of specialized datasets [26, 13, 25].
-Self-Instruct [36] is one popular method for generating synthetic datasets using an LLM.
-Following this methodology, Alpaca [32] generates 52K instruction-following demonstrations and subsequently fine-tunes the LLaMA 7B model [34] to create Alpaca 7B, which can behave qualitatively similarly to OpenAI’s text-davinci-003.
-Code Alpaca [5] extends this approach to generate code instructions using 21 seed tasks, while Gorilla-LM [29] adapts the method to focus on ML domain-specific APIs from Huggingface, TensorFlow Hub, and Torch Hub.
-To create more complex instructions, Evol-Instruct [40, 22] proposes iteratively updating instructions to become more complex through different prompting strategies.
-In addition to Evol-Instruct, OSS-Instruct [38] uses open-source code snippets to generate 75K high-quality instruction data and fine-tunes the Codelllama-Python-7B model to create Magicoder, which can match the performance of GPT-3.5-Turbo [27] on HumanEval [6].
-While these works focus on creating seed instruction sets to generate synthetic data for effectively fine-tuning an LLM, our research investigates post-processing methods in addition to Self-Instruct.
-Specifically, we concentrate on generating domain-specific programs in robotics [10], where we can effectively leverage constraints to filter out erroneous programs.
+To enhance LLMs’ performance in code generation, numerous studies have explored the creation of specialized datasets (Muennighoff et al., 2024; Köpf et al., 2023; Muennighoff et al., 2022).
+Self-Instruct (Wang et al., 2022) is one popular method for generating synthetic datasets using an LLM.
+Following this methodology, Alpaca (Taori et al., 2023) generates 52K instruction-following demonstrations and subsequently fine-tunes the LLaMA 7B model (Touvron et al., 2023) to create Alpaca 7B, which can behave qualitatively similarly to OpenAI’s text-davinci-003.
+Code Alpaca (Chaudhary, 2023) extends this approach to generate code instructions using 21 seed tasks, while Gorilla-LM (Patil et al., 2023) adapts the method to focus on ML domain-specific APIs from Huggingface, TensorFlow Hub, and Torch Hub.
+To create more complex instructions, Evol-Instruct (Xu et al., 2024; Luo et al., 2024) proposes iteratively updating instructions to become more complex through different prompting strategies.
+In addition to Evol-Instruct, OSS-Instruct (Wei et al., 2023) uses open-source code snippets to generate 75K high-quality instruction data and fine-tunes the CodeLlama-Python-7B model to create Magicoder, which can match the performance of GPT-3.5-Turbo (OpenAI, 2022) on HumanEval (Chen et al., 2021).
+While these works focus on creating seed instruction sets to generate synthetic data to effectively fine-tune an LLM, our research investigates post-processing methods in addition to Self-Instruct.
+Specifically, we concentrate on generating domain-specific programs in robotics (Hu et al., 2024), where we can effectively leverage constraints to filter out erroneous programs.
 6 Conclusion, Limitation and Future Works
 In this work, we introduce Robo-Instruct, a novel framework to generate synthetic training data to fine-tune small language models for domain-specific robot programs.
-Robo-Instruct comprises two novel components: 1) RoboSim, an angelic-execution-based algorithm to effectively validate Self-Instruct-generated programs, and 2) InstAlign, an instruction alignment procedure to revise instructions to better align with the generated programs.
-The experimental results demonstrate that the Codellama-Python-7B model fine-tuned on the Robo-Instruct-generated dataset can significantly outperform many popular open-weight LLMs for generating domain-specific robot programs.
-It also outperforms two proprietary LLMs, GPT-3.5-Turbo and Gemino-1.0-Pro, as well as the Self-Instruct-fine-tuned variant.
-A limitation of this study is that Robo-Instruct relies on Self-Instruct to filter invalid programs, making the dataset quality dependent on Self-Instruct’s performance.
-This can introduce biases if Self-Instruct consistently fails in certain areas.
-Future work will explore integrating Robo-Instruct with advanced methods like Evol-Inst and OSS-Inst to enhance dataset quality for domain-specific robot programs.
+Robo-Instruct comprises two novel components: 1) RoboSim, a method to synthesize consistent simulation environments dynamically for verifying the generated programs, and 2) InstAlign, an LLM-aided instruction alignment procedure to revise instructions to better align with the generated programs.
+The experimental results show that the 7B CodeLlama model fine-tuned on the Robo-Instruct dataset significantly outperforms larger open-weight LLMs and proprietary models like GPT-3.5-Turbo and Gemini-Pro in generating service robot programs.
+However, a key limitation is that Robo-Instruct relies on Self-Instruct to filter invalid programs, which may introduce biases and affect dataset quality.
+Another limitation is the use of a simple rejection-sampling method to handle invalid programs, which may not fully address the underlying issues.
+Future work will focus on improving dataset quality by integrating Robo-Instruct with advanced methods like Evol-Inst and OSS-Inst.
 Appendix A Appendix
 A.1 Overview
-In this appendix, we first present ablation experiments to investigate the percentage of invalid programs generated by Self-Instruct and examine how the generation temperature in InstAlign affects final performance.
-Next, we analyze and compare the datasets generated by Robo-Instruct and Self-Instruct.
-Finally, we list the seed tasks used in RoboEvaland the CoT prompt.
-A.2 Ablation Exmperiments
-Figure 5: Ablation Experiments
-A.2.1 effectivenss of the simulator
+In this appendix, we first outline the relationship between DynamicEval and the classic STRIPS planning formulation in subsection A.2, providing a new perspective on the proposed algorithm.
+In subsection A.3, we present additional ablation experiments to analyze the percentage of invalid programs generated by Self-Instruct and the effectiveness of the rejection-sampling strategy combined with RoboSim.
+We also explore how the generation temperature in InstAlign impacts final performance and compare the dataset diversity produced by Robo-Instruct and Self-Instruct.
+Subsection A.4 lists the seed tasks used in RoboEval and the CoT prompts.
+in subsection A.5, we report real-world experiments that empirically evaluate the performance of our fine-tuned model on two long-horizon tasks, which differ significantly from those in RoboEval, and assess the model’s latency in generating programs.
+Although this work focuses on service mobile robots, the proposed framework is adaptable to other domains.
+In subsection A.6, we offer toy examples showing how DynamicEval can be extended to verify programs by incorporating domain-specific constraints.
+A.2 Relevance to STRIPS planning
+The proposed DynamicEval shares significant similarities with the formulation of STRIPS planning.
+A STRIPS instance is typically represented as a tuple $\langle I,G,A,P\rangle$, where $I$ denotes the initial state of the simulation environment, $G$ represents the desired goal state that the robot aims to achieve, $A$ defines the set of actions available to transition between states, and $P$ is the set of preconditions that must be satisfied before performing actions.
+Thus, DynamicEval can be reformulated to align with the STRIPS formulation as shown in Alg.
+3.
+Each API invocation corresponds to an action, and its precondition consists of a set of literals, representing specific combinations of entities, types, and states.
+To address this, we extend the classic STRIPS formulation by incorporating dynamically discovered literals.
+Unlike the conventional STRIPS approach, where each literal is binary—True when defined and False when not—we introduce a third value, ”Undefined.” This means a literal must be explicitly defined as either True or False; otherwise, it remains in the Undefined state.
+When an action requires a literal that is undefined, a random value (True or False) is assigned to it, and the literal is added to the state of the simulation environment (line 7).
+Once the precondition is fully defined, the action is executed, and domain-specific constraints are checked for any violations (line 10).
+This extension enables DynamicEval to handle arbitrary programs effectively.
+[…]
+1:  Input: api_fn $\triangleright$ The API function name
+2:  Input: api_inputs $\triangleright$ The input received by the API invocation
+3:  Input: $\mathcal{W}$triangleright The current state of the simulation environment
+[…]
+5:  for $l\in p$ do $\triangleright$ Loop through every literal in the precondition
+6:   if $\textsc{CheckDefined}(\mathcal{W},l)\textbf{ is }\text{Undefined}$ then
+7:    $W\leftarrow\textsc{GrowWorld}(l,\mathcal{W})$triangleright Randonly instantiate the literal and grow $\mathcal{W}$ to include it
+8:   end if
+9:  end for
+10:  $\text{retval},\mathcal{W}\leftarrow\textsc{ExecUpdate}(\text{api\_fn},\text{params},\mathcal{W})$triangleright Execute api_fn and update $\mathcal{W}$
+11:  return  retval
+A.3 Ablation Exmperiments
+Figure 8: Ablation Experiment Results
+A.3.1 the Effectiveness of the Rejection-Sampling Strategy
 We analyze the percentage of instruction-program pairs discarded by RoboSim at various maximum resampling limits, as shown in Fig.
-5.
+8.
 Initially, with the maximum resampling limit set to 0, disabling the rejection-sampling method, approximately 51% of the programs generated by Self-Instruct contain errors.
@@
 However, there is a diminishing return; even with the maximum resampling limit set to 10, about 15% of the instructions still result in invalid programs.
-A.2.2 Instruction Alignment model temperature
-We further investigate how varying LLM temperatures for generating $\mathcal{I}_{\text{revised}}$ in InstAlign impact the performance of the fine-tuned model.
+A.3.2 Instruction Alignment model temperature
+We further investigate how varying LLM temperatures for generating the revised instruction in InstAlign impact the performance of the fine-tuned model.
 Fig.
@@
 As the temperature increases, we observe a decrease in performance.
-A.3 Analysis of the Generated Datasets
-(a) Token Length Distribution for Self-Instruct vs.Robo-Instruct
-(b) Cosine Similarity with RoboEvalfor Self-Instruct vs.Robo-Instruct
-Figure 6: Dataset Analysis
+A.3.3 Analysis of Generated Dataset
+Figure 9: Cosine similarities between RoboEval and generated data.
+Figure 10: Token length distribution for each instruction-program pair.
+Similar to Magicoder [Wei et al., 2023], we show the improvements from Robo-Instruct are not merely due to selection bias, i.e., including data more aligned with the distribution of RoboEval tasks than Self-Instruct.
+We pair each sample from the generated dataset with task instructions and their canonical solutions, then compute cosine similarity using TF-IDF embeddings [Sparck Jones, 1988].
+Fig.
+10 shows comparable token similarities between both methods.
+Additionally, Fig.
+10 presents the token length distribution, which also appears similar for both.
 Method Size Ngram=4 Score # Synth. Loc. # Synth. Obj.
+Self-Instruct 5K 0.581 956 1060
 Robo-Instruct 5K 0.587 1025 928
-Self-Instruct 5K 0.581 956 1060
 Table 4: Dataset Statistics
-We first compute and plot the distribution of token lengths in the Self-Instruct-generated dataset and the Robo-Instruct-generated dataset, as shown in Fig.
-6(a).
-Next, we measure the cosine similarity between each dataset and the RoboEvalbenchmark tasks following the approach in Magicoder [38], as depicted in Fig.
-6(b).
-Finally, Tab.
-3 presents the n-gram diversity score of each dataset, along with the number of synthesized locations and objects.
-Our findings indicate that both distributions and dataset statistics are very similar, suggesting that Robo-Instruct enhances the quality of the generated data over Self-Instruct rather than merely aligning the dataset towards the benchmark tasks.
-A.4 RoboEvalSeed Task Example
-1# Instruction: Go to Arjun’s office,
-2# ask him if he is ready to head out,
-3# and come back and tell me what he said
-4def task_program():
+Since RoboSim does not rely on pre-defined simulation environments, we aim to assess the diversity of programs generated by Self-Instruct and whether Robo-Instruct can maintain this diversity.
+To do so, we measure the number of distinct entities, such as synthetic locations and objects.
+As shown in Tab.
+4, with a dataset of only 5,000 samples, approximately 1,000 unique objects and locations are generated, highlighting that conventional robot simulations with pre-defined environments are insufficient.
+Additionally, Tab.
+4 presents the n-gram diversity scores for each dataset, indicating that both distributions and dataset statistics are highly similar.
+This suggests that Robo-Instruct not only preserves but enhances the quality of generated data compared to Self-Instruct, rather than simply aligning the dataset with benchmark tasks.
+A.4 Prompts
+A.4.1 RoboEval Seed Task Example
+Seed Task Example 1:
+1# Instruction: Go to Arjun’s office, ask him if he is ready to head out, and come back and tell me what he said
+2def task_program():
 3 start_loc = get_current_location()
@@
 9 say("Arjun said: " + response)
-Listing 1: Seed Task Example 1
+Seed Task Example 2:
 1# Instruction: Ask Alice if she needs 1, 2, or 3 boxes.
-2# Go to the storage room and ask if they have that many boxes.
-3# If so, go place the boxes in Alice’s office.
-4# Otherwise, tell Alice you could not get the boxes.
-5def task_program():
+Go to the storage room and ask if they have that many boxes.
+If so, go place the boxes in Alice’s office.
+Otherwise, tell Alice you could not get the boxes.
+2def task_program():
 3 go_to("Alice’s office")
@@
 19 say("I could not get the boxes")
-Listing 2: Seed Task Example 2
-1# Instruction: Check if there is a red marker in the main
-2# office, and if so, tell Eve that there is a marker there.
-3# If not, go to the supply room and
-4# bring a red marker to the main office.
-5def task_program():
+Seed Task Example 3:
+1# Instruction: Check if there is a red marker in the main office, and if so, tell Eve that there is a marker there. If not, go to the supply room and bring a red marker to the main office.
+2def task_program():
 3 go_to("main office")
@@
 12 place("red marker")
-Listing 3: Seed Task Example 3
-1# Instruction: Check every classroom if there is a whiteboard.
-2# Go to Aiden’s office to tell him which room does not
-3# have a whiteboard. Come back and tell me task is completed.
-4def task_program():
+Seed Task Example 4:
+1# Instruction: Check every classroom if there is a whiteboard. Go to Aiden’s office to tell him which room does not have a whiteboard. Come back and tell me task is completed.
+2def task_program():
 3 start_loc = get_current_location()
@@
 22 say("task is completed")
-Listing 4: Seed Task Example 4
-1# Instruction: Go to the kitchen and wait for someone
-2# to show up. When someone shows up, ask them to open
-3# the fridge, then pick up a diet coke.
-4# Finally, put the diet coke in the living room.
-5def task_program():
+Seed Task Example 5:
+1# Instruction: Go to the kitchen and wait for someone to show up. When someone shows up, ask them to open the fridge, then pick up a diet coke. Finally, put the diet coke in the living room.
+2def task_program():
 3 go_to("kitchen")
@@
 14 place("diet coke")
-Listing 5: Seed Task Example 5
-1# Instruction: Take a bed sheet from the laundry room
-2# and put it in each of the bedrooms.
-3def task_program():
+Seed Task Example 6:
+1# Instruction: Take a bed sheet from the laundry room and put it in each of the bedrooms.
+2def task_program():
 3 start_loc = get_current_location()
@@
 12 go_to(start_loc)
-Listing 6: Seed Task Example 6
-A.5 Prompts to Generate Synthetic Dataset Using Self-Instruct
-You are a helpful assistant. Here is a robot that has the following capabilities:
-- def get_current_location() -> str:
-- def get_all_rooms() -> list[str]:
-- def is_in_room(object : str) -> bool:
-- def go_to(location : str) -> None:
-- def ask(person : str, question : str, options: list[str]) -> str:
-- def say(message : str) -> None:
-- def pick(obj: str) -> None:
-- def place(obj: str) -> None:
-Generate an interesting robot task that can be accomplished using the above capabilities.
-{{SEED EXAMPLE}}
-Generate an interesting robot task that can be accomplished using the above capabilities.
-Table 4: Prompts to Generate Synthetic Dataset Using Self-Instruct.
-A.6 CoT Prompts for InstAlign
-### Role: You are an expert at understanding robot programs.
-You will be given a task instruction and robot program pair.
-However, the instruction may not align with the program well.
-You need to correct the task instruction to match the given robot program.
-### Context: The robot only has access to the following 8 APIs and standard Python functions
-- def get_current_location() -> str:
-- def get_all_rooms() -> list[str]:
-- def is_in_room(object : str) -> bool:
-- def go_to(location : str) -> None:
-- ask(person : str, question : str, options: list[str]) -> str:
-- say(message : str) -> None:
-- def pick(obj: str) -> None:
-- def place(obj: str) -> None:
-### Inputs
-Original Instruction: This is a task instruction that may not align with the robot program Robot Program: This is a python function starting with ‘def task_program():‘
-### Task:
-1. Write down all the provided APIs used in the program and explain the effect of each API in this program
+A.4.2 Prompts to Generate Synthetic Dataset Using Self-Instruct
+1You are a helpful assistant. Here is a robot that has the
+2following capabilities:
+3- def get_current_location() -> str:
+4- def get_all_rooms() -> list[str]:
+5- def is_in_room(object : str) -> bool:
+6- def go_to(location : str) -> None:
+7- def ask(person : str, question : str, options: list[str]) -> str:
+8- def say(message : str) -> None:
+9- def pick(obj: str) -> None:
+10- def place(obj: str) -> None:
+11Generate an interesting robot task that can be accomplished using the
+12above capabilities.
+13{SEED EXAMPLE 1}
+14
+15...
+16
+17Generate an interesting robot task that can be accomplished using the
+18above capabilities.
+19{SEED EXAMPLE 6}
+20
+21Generate an interesting robot task that can be accomplished using the
+22above capabilities.
+A.4.3 CoT Prompts for InstAlign
+1### Role
+2You are an expert at understanding robot programs.
+3You will be given a task instruction and robot program pair.
+4However, the instruction may not align with the program well.
+5You need to correct the task instruction to match the given robot program.
+6
+7### Context
+8The robot only has access to the following 8 APIs and
+9standard Python functions
+10- def get_current_location() -> str:
+11- def get_all_rooms() -> list[str]:
+12- def is_in_room(object : str) -> bool:
+13- def go_to(location : str) -> None:
+14- ask(person : str, question : str, options: list[str]) -> str:
+15- say(message : str) -> None:
+16- def pick(obj: str) -> None:
+17- def place(obj: str) -> None:
+18
+19### Inputs
+20Original Instruction
+21 This is a task instruction that may not align with the robot program
+22Robot Program
+23 This is a python function starting with def task_program():‘
+24
+25### Task
+261. Write down all the provided APIs used in the program and
+27explain the effect of each API in this program
 282. Examine these APIs and write down step by step what the program does
-3. Combine all the results above and rewrite the instruction under # Final Corrected Instruction: You need to be specific and clear in your final corrected instruction.
-Table 5: CoT Prompts for InstAlign.
+293. Combine all the results above and rewrite the instruction
+30You need to be specific and clear in your final corrected instruction.
+A.5 Real-world Experiments
+A.5.1 Program Generation Results Beyond RoboEval
+We conducted a small qualitative experiment to assess the performance of the base model, Self-Instruct fine-tuned models, Robo-Instruct fine-tuned models, and GPT-4 on long-horizon tasks.
+To showcase performance, we created two new long-horizon tasks that differ significantly from those in the RoboEval benchmark.
+Using a temperature setting of 0, we generated programs with each model and found that our Robo-Instruct fine-tuned model successfully produced correct programs, while both the base model and Self-Instruct fine-tuned model failed.
+Additionally, GPT-4 made an error on the second instruction.
+We show the generation results from Robo-Instruct fine-tuned models and GPT-4 below.
+Task instruction 1:
+1Lets play a game: Double and give it to the next person.
+Start with 1 dollar.
+Go to rooms A, B, C, D, E, F, and G.
+If you see someone, tell them how much money you have.
+Then ask if they would like to take the money now or double the amount and give it to the next person.
+If they choose to take it, the game is over, and you should come back to me.
+Otherwise, double your money and continue.
+If, in the end, no one takes the money, tell me how much you still have.
+Robo-Instruct fine-tuned model (Ours) [Success]
+1def task_program():
+2 start_loc = get_current_location()
+3 list_of_rooms = ["A", "B", "C", "D", "E", "F", "G"]
+4 money = 1
+5 for room in list_of_rooms:
+6 go_to(room)
+7 if is_in_room("person"):
+8 say("I have \$" + str(money))
+9 response = ask("", "Would you like to take the money now
+10 or double the amount and give it to the next person?",
+11 ["Take now", "Double and give"])
+12 if response == "Take now":
+13 go_to(start_loc)
+14 say("The game is over, I gave the money to " + room)
+15 return
+16 else:
+17 money *= 2
+18 go_to(start_loc)
+19 say("No one took the money, I still have \$" + str(money))
+GPT-4 [Success]
+1def task_program():
+2 start_loc = get_current_location()
+3 rooms = ["A", "B", "C", "D", "E", "F", "G"]
+4 money = 1
+5 for room in rooms:
+6 go_to(room)
+7 if is_in_room("person"):
+8 say("I have " + str(money) + " dollars")
+9 response = ask("", "Would you like to take the money now
+10 or double the amount and give it to the next person?",
+11 ["Take it", "Double it"])
+12 if response == "Take it":
+13 say("Game over. The person in room " + room + " took the money.")
+14 go_to(start_loc)
+15 return
+16 else:
+17 money *= 2
+18 go_to(start_loc)
+19 say("No one took the money. I still have " + str(money) + " dollars.")
+Task instruction 2:
+1Go to my office and check if I have a table, a chair, and a monitor there.
+If any of these items are missing, go to Jasons office and see if he is there.
+If he is, ask him if I can borrow the missing items.
+If he agrees, pick up each missing item and bring it to my office.
+If Jason is not in his office or he says no, come back and tell me the
+2reason.
+Robo-Instruct fine-tuned model (Ours) [Success]
+1def task_program():
+2 start_loc = get_current_location()
+3 go_to("my office")
+4 table_found = is_in_room("table")
+5 chair_found = is_in_room("chair")
+6 monitor_found = is_in_room("monitor")
+7 if not table_found:
+8 go_to("Jason’s office")
+9 if is_in_room("Jason"):
+10 response = ask("Jason", "Can I borrow the table from you?",
+11 ["Yes", "No"])
+12 if response == "Yes":
+13 go_to("Jason’s office")
+14 pick("table")
+15 go_to("my office")
+16 place("table")
+17 else:
+18 go_to(start_loc)
+19 say("Jason said no")
+20 else:
+21 go_to(start_loc)
+22 say("Jason is not in his office")
+23 if not chair_found:
+24 go_to("Jason’s office")
+25 if is_in_room("Jason"):
+26 response = ask("Jason", "Can I borrow the chair from you?",
+27 ["Yes", "No"])
+28 if response == "Yes":
+29 go_to("Jason’s office")
+30 pick("chair")
+31 go_to("my office")
+32 place("chair")
+33 else:
+34 go_to(start_loc)
+35 say("Jason said no")
+36 else:
+37 go_to(start_loc)
+38 say("Jason is not in his office")
+39 if not monitor_found:
+40 go_to("Jason’s office")
+41 if is_in_room("Jason"):
+42 response = ask("Jason", "Can I borrow the monitor from you?",
+43 ["Yes", "No"])
+44 if response == "Yes":
+45 go_to("Jason’s office")
+46 pick("monitor")
+47 go_to("my office")
+48 place("monitor")
+49 else:
+50 go_to(start_loc)
+51 say("Jason said no")
+52 else:
+53 go_to(start_loc)
+54 say("Jason is not in his office")
+55 go_to(start_loc)
+56 say("task is completed")
+GPT-4 [Failure]
+1def task_program():
+2 start_loc = get_current_location()
+3 go_to("my office")
+4 missing_items = []
+5 for item in ["table", "chair", "monitor"]:
+6 if not is_in_room(item):
+7 missing_items.append(item)
+8 if len(missing_items) > 0:
+9 go_to("Jason’s office")
+10 if is_in_room("Jason"):
+11 for item in missing_items:
+12 response = ask("Jason", "Can I borrow your " + item +
+13 "?", ["Yes", "No"])
+14 if response == "Yes":
+15 # BUG: did not return to Jason’s office after going back to my office
+16 pick(item)
+17 go_to("my office")
+18 place(item)
+19 else:
+20 go_to(start_loc)
+21 say("Jason did not allow me to borrow his " + item)
+22 return
+23 else:
+24 go_to(start_loc)
+25 say("Jason is not in his office")
+26 return
+27 go_to(start_loc)
+28 say("Task is completed")
+A.5.2 Generation Latency
+(a) Example Program 1 generated by our fine-tuned model.
+(b) Example Program 1 generated by GPT-4.
+Figure 11: Our fine-tuned model is approximately 6x faster in inference speed than GPT-4 (Part 1).
+(c) Another Example Program generated by our fine-tuned model.
+(d) Another Example Program generated by GPT-4.
+Figure 12: Our fine-tuned model is approximately 6x faster in inference speed than GPT-4 (Part 2).
+A.6 Toy Examples Beyond Service Mobile Robots
+A.6.1 Robot with low-level controls
+Consider a tabletop manipulation scenario with a potential API function, is_rotate(robot_gripper_ name, radians), where the robot’s gripper has a physical constraint, allowing rotation only within the range $[-\frac{\pi}{6},\frac{\pi}{6}]$ radians.
+For the following generated program snippet:
+1def task_program():
+2 rotate("left hand", math.pi/6)
+3 rotate("left hand", math.pi/6)
+4 rotate("left hand", math.pi/6)
+DynamicEval will first infer that ”left hand” is an entity of the robot gripper type.
+Then, regardless of the initial configuration of the gripper, DynamicEval will throw an error because the program causes the gripper to exceed its allowable range of motion.
+A.6.2 AI-powered personal digital assistant
+Consider a broader application than robotics: code generation for an AI-powered personal digital assistant.
+This AI assistant could handle scheduling events using an API function like schedule_on_calendar(event, start_time, duration).
+Given the instruction: ”My schedule is free tomorrow morning.
+Please create two 1-hour timeslots for office hours for my robotics and deep learning class.” The assistant could generate a program to create these timeslots:
+1def task_program():
+2 schedule_on_calendar("robotics class office hour",
+3 "9:30 am", "1 hr")
+4 schedule_on_calendar("deep learning class office hour",
+5 "10:00 am", "1 hr")
+In this example, DynamicEval needs to reason about the entities “robotics class office hour” and “deep learning class office hour”, which are categorized as event types.
+The event type indicates that these entities have associated timeslots.
+The state of these entities is defined by the time they occur: robotics class office hour is set for 9:30-10:30 am, and deep learning class office hour is set for 10:00-11:00 am.
+During evaluation, DynamicEval can identify a time conflict between these two office hours and thus determine that the generated program is invalid.